OpenStudy (anonymous):

Using the domain values {-1, 0, 4}, find the range values for relation f(x)=3x-8. Range={-11, -8, 4} Range={-11, 8, 12} Range={-11, -8, -4} None of the above

OpenStudy (anonymous):

how can i do that cause its next to f.. do i classify f as 1 ?

OpenStudy (anonymous):

Put where x -1, 0 and 4 so you take x= -1, y= -9 x=0, y= -7 x=4, y=1 so Range={-9, -7, 1}

OpenStudy (anonymous):

oh .

OpenStudy (anonymous):

same thing..substitute x=-1 ...find y... x=0 find y ....x=4 find y.!!

OpenStudy (anonymous):

i got confused for a minute.. sorry!

OpenStudy (anonymous):

the first option is correct

OpenStudy (anonymous):

\[\huge none \quad of \quad the \quad ABOVE\]

OpenStudy (anonymous):

PROOF @sauravshakya

OpenStudy (anonymous):

can you help me with my next question please

OpenStudy (anonymous):

thank u rebecca

OpenStudy (anonymous):

okay msg me the url

OpenStudy (anonymous):

when x = -1 then f(-1)=3*(-1)-8=-11 when x = 0 then f0) = 3*0-8=-8 when x = 3 then f(4)= 3*4-8=4 so range = {-11,-8,4)